ABC Company purchased a new machinery two years ago for $68759. Today, it is selling this machinery for $25416. What is the after-tax salvage value if the tax rate is 37 percent?

The MACRS allowance percentages are as follows, commencing with year one: 20.00, 32.00, 19.20, 11.52, 11.52, and 5.76 percent.
Answer #1

Book value as on date of sale=cost-accumulated depreciation

=68759(1-0.2-0.32)

=$33004.32

Hence loss on sales =33004.32-25416

=7588.32

Hence after tax salvage value=sale proceeds+(tax rate*loss on sale)

=25416+(7588.32*0.37)

=$28223.6784